You are at the buger barn. Two hamburgers and one hot dog cost $4.82. At the same price, one hamburger and two hotdogs cost $3.70. How much does a hotdog cost?
Answer by stanbon(48558)

Let x be cost of a hamburger and y be cost of hot dog
2x+y=482 cents
x+2y=370 cents
Multiply the 2nd eq. by "2" to get 2x+47=740

Subtract this from the 1st eq. to get:
-3y=-258
y=86 cents
Substitute that into the 1st equation to get:
2x+86=482
2x=396
x=198 cents
So, hambugers cost $1.98; hot dogs cost $.86
Cheers,
Stan H.